Please show WORK!

In one species of rabbits, fur color is (hypothetically) controlled by a single gene, with dark fur being dominant to light fur. You study a wild population of these rabbits, and determine that 84% have dark fur.

1. If the population is in Hardy-Weinberg equillibrium, what would be the percentages of each genotype? (Homozygous dominant, Heterozygous, and Homozygous recessive)

2. The following spring, you examine 500 newly born babies, and observe that 390 of these have dark fur. Perform a Chi-Squared test (use p=0.05) to determine whether this generation remains in Hardy-Weinberg equillibrium. Be sure to make a conclusion based on these results (and phenotypes that you observe).

- What is your Chi-Squared value

- How many degrees of freedom

- What is your critical value (p=0.05)

- Is this population still in Hardy-Weinberg equillibrium? Why or why not?

3. If the population had 84% dark fur and had significant inbreeding (F= 0.2), what percentage of the next generation would you have expected to be heterozygous?
Answer #1

The Phenotype expressing dominant D is 84% of total population.

The phenotype expressing recessive d is the remaining population i.e 16%.(100-84)

The phenotype expressing dominant D is the sum of homozygous dominant (DD) and heterozygous dominant (Dd) population.

for easy calculation lets assume population of 100.if this population is in hardy weinberg equilbrium

(D+d)2 = 100

DD+2Dd= 84 (84% *100) (given)

(D+d)2 = D2+2Dd+d2 =100

84+d2 =100

d=4

(D+d)2 =100

(D+d)=10

Sub d=4 we get D= 6

So in a population in Hardy weinberg equilbrium

The Homozygote Dominant is = D2 = 36

The Heterozygote Dominant = 2Dd=48

Homozygote Dominant = d2=16

Since total population is 100 we can consider the above values in percentage itself (36/100*100)

2) Among 500 new born, 390 of them express dominant phenotype (D)

For a population in Hardy weinberg equilbrium phenotypes expressing dominant phenotype is 84 % of total population.In this case total population is 500 .so the expected number= (84/100 *500) = 420

To understand whether the population is following Hardy weinberg ,

we have to test the goodness of fit (Χ2) for the expected (E) and observed (O) population

Χ2 = (O-E)2 / E = (390-420)2/ 420 = 2.14

if the expected and observed values are equal the Χ2= 0 ( Note: The probability of this happening for all practical purpose is zero or p=0)

So we assume a significant p value to see the closeness of two population (p=0.05,this is accepted level of significance for large population)

Now we need to know the Χ2 value for p=0.05.This value is found in the Χ2 table . that is given below .We can see that we need to find a value degree of freedom to get the Χ2 value at p= 0.05 .

The degree of freedom is nothing but value  = no of expected categories -1 .The no of expected categories or genotypes is  2 (DD or Dd) .So degree of freedom =1(2-1)

Chi-Square (?) Distribution Area to the Right of Critical Value 0.95 0.90 0.10 0.05 Degrees of Freedom 0.99 0.975 0.025 0.01

From the above table Χ2 value with degree of freedom (1) and at (p=0.05)

Χ2 = 3.841

The two are same only if calculated Χ2 value < Χ2 value at p=0.05 or Χ2 Critical value

the calculated Χ2 value is 2.14 and  Critical Χ2 value is 3.84

Since it is less than critical value , the two population are similar and follow Hardy Weinberg Equilbrium

3)

Inbreeding always causes a decrease in heterozygote population This is computed from the F coefficent for inbreeding.

Hin = Hwb (1-F)

For the population in hardy weinberg equilbrium Heterozygote(Dd) population is 48 (Calculation shown in 1st problem)

Hin = 48 (1-0.2) =38.4

38.4 % of the next generation of population (inbreeding coefficent =0.2) will be heterozygotes.